OpenStudy (anonymous):

the everton college store paid $1523 for and order of 45 calculators. The store paid $9 for each scientific calculator. The others, all graphing calculators, cost the store $52 each how many of each type of calculator was ordered?

OpenStudy (anonymous):

store purchased 19 scientific calculators and 26 graphing calculators If u need the working, let me know

OpenStudy (anonymous):

let x=$9 let y=$52 x+y=45 y=45-x 9x+52y=$1523 Subtitution y=45-x 9x+52(45-x)=$1523 9x+2,340-52x=$1,523 -43x =$1,523-2,340 x=-817/-43 x = 19 Subtitution x = 19 solve for y x+y=45 19+y=45 y=45-19 y =26

OpenStudy (anonymous):

lick on god answer if I deserve

OpenStudy (anonymous):

@Nancy Had u taken one as x and the other as 45-x then calculation wud be shorter and much simpler ( :}

OpenStudy (anonymous):

thank harkirat,I try sept by step explant

OpenStudy (anonymous):

I know and appreciate, but what i suggest wud be much simpler, try n see...

OpenStudy (anonymous):

I know short fast ,but I try made undestand by step by step